OpenStudy (liziekay):

Find the remaining trigonometric ratios of θ based on the given information. csc θ = 2 and cos θ is negative sin θ = 1/2 cos θ = tan θ = cot θ = sec θ =

OpenStudy (evoker):

Basically you need to figure out y,x, and h, and then use the trig relationships to answer all the questions, just as a reminder to get started sin=y/h

OpenStudy (liziekay):

How would we find the hypotenuse?

OpenStudy (liziekay):

I mean the x? Adjacent?

OpenStudy (liziekay):

@Evoker

OpenStudy (visarika23):

Cos(theta) is negative which means theta may be in Q2 orQ3 Now Sin(theta) is positive so theta must be in Q2 In Q2 only sin and cosec are positive and all other trigonometric ratios are negative Therefore put the sign accordingly

OpenStudy (liziekay):

would the adjacent x be the square root of 3? @Visarika23

OpenStudy (visarika23):

Yup Sin(theta) =1/2=y/h x^2=h^2-y^2=4-1=3 x=square root 3

OpenStudy (visarika23):

Now u know the values of x,y,h. So u can find out the values of all trigonometric ratios with their signs

OpenStudy (liziekay):

Great! Thanks

OpenStudy (visarika23):

Yr wlcm